STUDYING CHEMISTRY WITH VIBRATIONAL OVERTONE PHOTODISSOCIATION SPECTROSCOPY

\underline{SERGEY A NIZKORODOV}, Department of Chemistry, University of California at Irvine, Irvine CA 92697-2025.

This talk will discuss applications of vibrational overtone photodissociation spectroscopy to problems in molecular reaction dynamics and chemistry of weakly-bound molecules. Specific examples from the past research experience of the presenter will include: i) vibrationally-mediated photodissociation of water dimer and Argon-water complexes; ii) predissociation spectroscopy of HOONO isomer of nitric acid; iii) overtone photodissociation of peroxynitric acid (PNA).